OpenStudy (anonymous):

Solve –3(2x – 4) = 24

OpenStudy (anonymous):

-6x+12=24 ==> x=-2

OpenStudy (anonymous):

\[-3(2x)-3(-4) = -6x+12\] \[-6x+12=24\] \[-6x=24-12\] \[-6x=12\] \[x=\frac{12}{-6}\] \[x=-2\]